OpenStudy (anonymous):

Simplify. (2x((cubed))y(tiny 5))) (tiny 4) A. 16x(tiny 7)y(tiny 9) B. 8x(tiny 12)y(tiny 20) C. 2x(tiny 7)y(tiny 9) D. 16x(tiny 12)y(tiny 20)

OpenStudy (anonymous):

That's what it looks like, in the attached file

OpenStudy (anonymous):

@Igreen

OpenStudy (anonymous):

@gabbyalicorn

OpenStudy (alyssa_xo):

2 properties to remember. \((ab)^M=a^M \times b^M\\ {(a^M)}^N=a^{M*N}\\\)

OpenStudy (anonymous):

IDK how to do this at all!